# Probabilistic Control Barrier Functions: Safety in Probability for Discrete-Time Stochastic Systems

### v1

@@ Abstract
Control systems operating in the real world face countless sources of unpredictable uncertainties. These random disturbances can render deterministic guarantees inapplicable and cause catastrophic safety failures. To overcome this, this paper proposes a method for designing safe controllers for discrete-time stochastic systems that retain probabilistic guarantees of safety. To do this we modify the traditional notion of a control barrier function (CBF) to explicitly account for these stochastic uncertainties and call these new modified functions probabilistic CBFs. We show that probabilistic CBFs can be used to design controllers that guarantee safety over a finite number of time steps with a prescribed probability. Next, by leveraging various uncertainty quantification methods, such as concentration inequalities, the scenario approach, and conformal prediction, we provide a variety of sufficient conditions that result in computationally tractable controllers with tunable probabilistic guarantees across a plethora of practical scenarios. Finally, we showcase the applicability of our results in simulation and hardware for the control of a quadruped robot.
@@ I-B Statement of Contributions
We consider the problem of designing safe controllers for discrete-time stochastic systems. Our first contribution is the introduction of a probabilistic notion of CBF which we show can be used to design controllers with probabilistic safety guarantees over a finite time horizon. However, verifying whether a function is a probabilistic CBF is impractical, because it requires knowledge of the uncertainty distribution. In our second contribution we sidestep this difficulty by introducing two different sufficient conditions which guarantee that a function is a probabilistic CBF. These conditions only require knowledge of the moments of the distribution of the value of the CBF at the next time step. The first condition only leverages a known upper bound on the CBF and the first moment of such distribution, whereas the second condition uses its first and second moments. We also study various cases under which these conditions are convex in the control input and are thus amenable to optimization-based control design. Next, we consider the problem of verifying that a function is a probabilistic CBF when moment information is not available but instead we have a dataset consisting of uncertainty realizations. In our third contribution we leverage concentration inequalities, scenario optimization, and conformal prediction to provide different sufficient conditions based on such dataset that ensure that a function is a probabilistic CBF with high confidence. Finally, we validate our results in both simulation and hardware on quadruped robots.
@@ IV Verification of Probabilistic CBFs Using Data
In the rest of the section, we fix $\delta,\beta\in(0,1)$, and consider a dataset ${\mathbf{D}}_{N}$ of $N\in\mathbb{N}$ samples as defined in Proposition 6. The following result establishes a sufficient condition for $h$ to be a probabilistic CBF with a given confidence in the case where the random variable $\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}}))$ is bounded.
@@
(Hoeffding-based condition): Suppose that there exist constants $a,b,\epsilon>0$ with $\beta\geq 2\exp\{-\frac{2N\epsilon^{2}}{(b-a)^{2}}\}$ and such that for each ${\mathbf{x}}\in{\mathcal{C}}$ and ${\mathbf{u}}\in\real^{m}$, $a\leq\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})\leq b$ almost surely and
@@
Let ${\mathbf{x}}\in{\mathcal{C}}$ and pick ${\mathbf{u}}\in\real^{m}$ so that (17) holds. Note that such $\mathbf{u}$} is dependent on the dataset ${\mathbf{D}}_{N}$. Now, we have
@@
$\geq 1-2\exp\Bigg\{-\frac{2N\epsilon^{2}}{(b-a)^{2}}\Bigg\}\geq 1-\beta,$
where in the first inequality we have used Hoeffding’s inequality [39], and in the last inequality we have used $\beta\geq 2\exp\{-\frac{2N\epsilon^{2}}{(b-a)^{2}}\}$. This, together with (17), implies that
@@
(Conditions of Proposition 7): A simple scenario in which $a,b$ as in Proposition 7 exist is if $h$ is uniformly bounded (which we can always assume without loss of generality). Moreover, $\beta\geq 2\exp\{-\frac{2N\epsilon^{2}}{(b-a)^{2}}\}$ holds if $N$ is sufficiently large. We note that we can also state a version of Proposition 7 in which the $a,b,\epsilon,N$ are dependent on $\mathbf{x}$} and $\mathbf{u}$} (instead of being uniform for all ${\mathbf{x}}\in{\mathcal{C}}$ and ${\mathbf{u}}\in\real^{m}$). Although the assumption that $a,b,\epsilon,N$ are uniform across $\mathbf{x}$} and $\mathbf{u}$} might add some conservatism, it leads to a simpler dependency on $\mathbf{u}$} in (17). In fact, condition (17) is convex in $\mathbf{u}$} under assumptions such as the ones in Remark 1.  $\bullet$
@@
which is feasible by assumption. Now, by [40, Theorem 2.4] we have that for each ${\mathbf{x}}\in{\mathcal{C}}$,
$\mathbb{P}_{N}\Big(\mathbb{P}(\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})\leq 0)>\delta\Big)\leq\beta.$
Equivalently,
@@
Another method to establish a distribution-free guarantee that $h$ is a probabilistic CBF with a given confidence is through the theory of conformal prediction [42]. This is leveraged in the following result.
Proposition 9.
(Conformal prediction-based condition): Suppose that for each ${\mathbf{x}}\in{\mathcal{C}}$, there exists ${\mathbf{u}}\in\real^{m}$ such that
$Q_{N}=\text{Quantile}_{1-\delta+\sqrt{\frac{\log(1/\beta)}{2N}}}\Big(R^{(1)},\ldots,R^{(N)},\infty\Big)\leq 0,$ (22)
with $R^{(i)}=-\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}}^{(i)})$. Then, $h$ is a $\delta$-probabilistic CBF with confidence $1-\beta$.
Proof.
Let ${\mathbf{x}}\in{\mathcal{C}}$ and pick ${\mathbf{u}}\in\real^{m}$ so that (22) holds. Note that such $\mathbf{u}$} is dependent on the dataset ${\mathbf{D}}_{N}$. By [42, Proposition 2a], we have
$\mathbb{P}_{N}\Big(\mathbb{P}(-\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})\leq Q_{N})\geq 1-\delta\Big)\geq 1-\beta.$ (23)
Note that if ${\mathbf{d}}\in\real^{d}$ satisfies $\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})\leq Q_{N}$, then we have $\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})\leq 0$ by (22). Hence, for any fixed dataset ${\mathbf{D}}_{N}$ and $\mathbf{u}$} satisfying (22), we have
$\mathbb{P}\Big(-\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})\leq Q_{N}\Big)\leq\mathbb{P}\Big(-\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})\leq 0\Big).$
Therefore, if the event $\bar{X}_{N}$ defined as $\mathbb{P}\Big(-\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})\leq Q_{N}\Big)\geq 1-\delta$ holds, then the event $\bar{Y}_{N}$ defined as $\mathbb{P}\Big(-\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})\leq 0\Big)\geq 1-\delta$ also holds. Hence, $\mathbb{P}_{N}(\bar{Y}_{N})\geq\mathbb{P}_{N}(\bar{X}_{N})$ and by (23) this implies that $h$ is a $\delta$-probabilistic CBF with confidence $1-\beta$. ∎
Since condition (22) is derived through conformal prediction, we refer to it as the conformal prediction-based condition. We note that as shown in [43, Section 5], the quantile constraint (22) can be reformulated using mixed-integer programming so that it can be used in optimization-based control design with off-the-shelf solvers.
@@ V-A Simulation
In simulation, we model the uncertainties in the terrain through a Gaussian disturbance with standard deviation $\sigma$ acting only in the $y$ direction, i.e., ${\mathbf{d}}_{k}=[0,d_{k},0]$, with $d_{k}\sim{\mathcal{N}}(0,\sigma)$. We consider a time horizon of 20 steps and take $\delta=0.1$, $\beta=0.01$, $\alpha=0.01$ and $\Delta t=0.1$. We consider a nominal controller ${\mathbf{k}}_{\text{nom}}({\mathbf{x}})=[0.2,0,-\theta]$. For each of the different conditions ——Markov (in the form of (8)), Cantelli, Hoeffding, scenario approach, and conformal prediction——, we compute the control input at every state through a safety filter, i.e., by solving an optimization problem that finds the closest input to the nominal controller that satisfies the corresponding condition. We generate datasets of size 3252, 113, and 300 for the Hoeffding, scenario, and conformal approaches, respectively. For the Markov, Hoeffding and scenario approach-based condition, the optimization problem at every state is a convex quadratically constrained quadratic program (QCQP), which we solve using the cvxpy library in Python. The Cantelli-based condition leads to a non-convex problem, which we solve using the trust region solver of the minimize function in the SciPy library in Python. Using the reformulation in [43, Section 5], the conformal prediction-based method leads to a mixed-integer quadratically constrained quadratic program, which we solve using the gurobipy library in Python. We implement each of the approaches (along with the nominal controller and the approach based on [27], which we refer to as the Martingale approach) in 400 different trajectories with initial condition ${\mathbf{x}}_{0}=[0,0,0]$.
@@
Figure 1: (left) Number of unsafe trajectories for different values of the disturbance variance $\sigma$. (right) Number of unsafe trajectories for different values of the trajectory length. No filter refers to the implementation of the nominal controller, whereas Markov, Cantelli, Hoeffding, Scenario, and Conformal refer to the approaches in Propositions 2789, respectively.
@@
(b) Cantelli
(c) Hoeffding
(d) Scenario
(e) Conformal
Figure 2: Evolution of 400 different trajectories with initial condition at ${\mathbf{x}}_{0}=[0,0,0]$ and $\sigma=0.06$ and trajectory length 2020 for different methods introduced in the paper, along with the nominal controller (left).
Markov Cantelli Hoeffding Scenario Conformal Martingale
$T_{ex}$ 3.47 3.78 3.89 79.96 5.65 3.46
$\sigma_{0}$ 0.03 0.16 0.13 0.21 0.2 0.25
TABLE I: Average time it takes to compute the controller ($T_{ex}$) in milliseconds and smallest value of $\sigma$ that leads to infeasibility $\sigma_{0}$ for the different approaches.
Finally, Table I shows the average execution times for the different approaches as well as the smallest value of $\sigma$ that leads to infeasibility of the corresponding condition.
The choice of which of the different approaches should be used depends on a variety of factors. Firstly, on what information is available about the uncertainty distribution. Secondly, on the desired level of conservativeness with respect to the safety constraints. Figure 1 and Table I show that there is a tradeoff between the empirical number of safety violations and the smallest variance that leads to infeasibility. Approaches yielding less safety violations generally lead to infeasible conditions for smaller values of $\sigma$. Thirdly, another point to consider is the computational time (cf. Table I). Although the martingale approach shows a larger value of $\sigma_{0}$ in Table I, it fails to meet the prescribed safety guarantees that the other approaches successfully satisfy.

### v2

@@ Abstract
Control systems operating in the real world face countless sources of unpredictable uncertainties. These random disturbances can render deterministic guarantees inapplicable and cause catastrophic safety failures. To overcome this, this paper proposes a method for designing safe controllers for discrete-time stochastic systems that retain probabilistic guarantees of safety. To do this we modify the traditional notion of a control barrier function (CBF) to explicitly account for these stochastic uncertainties and call these new modified functions probabilistic CBFs. We show that probabilistic CBFs can be used to design controllers that guarantee safety over a finite number of time steps with a prescribed probability. Next, by leveraging various uncertainty quantification methods, such as concentration inequalities and the scenario approach, we provide a variety of sufficient conditions that result in computationally tractable controllers with tunable probabilistic guarantees across a plethora of practical scenarios. Finally, we showcase the applicability of our results in simulation and hardware for the control of a quadruped robot.
@@ I-B Statement of Contributions
We consider the problem of designing safe controllers for discrete-time stochastic systems. Our first contribution is the introduction of a probabilistic notion of CBF which we show can be used to design controllers with probabilistic safety guarantees over a finite time horizon. However, verifying whether a function is a probabilistic CBF is impractical, because it requires knowledge of the uncertainty distribution. In our second contribution we sidestep this difficulty by introducing two different sufficient conditions which guarantee that a function is a probabilistic CBF. These conditions only require knowledge of the moments of the distribution of the value of the CBF at the next time step. The first condition only leverages a known upper bound on the CBF and the first moment of such distribution, whereas the second condition uses its first and second moments. We also study various cases under which these conditions are convex in the control input and are thus amenable to optimization-based control design. Next, we consider the problem of verifying that a function is a probabilistic CBF when moment information is not available but instead we have a dataset consisting of uncertainty realizations. In our third contribution we leverage concentration inequalities and scenario optimization to provide different sufficient conditions based on such dataset that ensure that a function is a probabilistic CBF with high confidence. Finally, we validate our results in both simulation and hardware on quadruped robots. Our main contributions are summarized in Table I.
Result Assumption Technique Guarantee
Prop. 2 Mean Markov’s inequality Single prob.
Prop. 4 Mean ++ variance Cantelli’s inequality Single prob.
Prop. 7 Samples Hoeffding’s inequality Double prob.
Prop. 8 Samples Scenario approach Double prob.
TABLE I: Table summarizing of the main results in the paper. The first column indicates the result, the second column the assumptions on the uncertainty needed to apply the result, the third column the main technical tool used in its derivation, and the fourth column indicates the type of guarantee it provides, which can be either single probability (as in Definition 2) or double probability (as in Definition 3).
@@ IV Verification of Probabilistic CBFs Using Data
In the rest of the section, we fix $\delta,\beta\in(0,1)$, and consider a dataset ${\mathbf{D}}_{N}$ of $N\in\mathbb{N}$ samples as defined in Proposition 6. The following result establishes a sufficient condition for $h$ to be a probabilistic CBF with a given confidence in the case where the random variable $\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})$ is bounded and has a special separable form.
@@
(Hoeffding-based condition): Suppose that there exist functions $H_{1},H_{3}:\real^{n}\times\real^{m}\to\real$, $H_{2},H_{4}:\real^{n}\times\real^{d}\to\real$ such that
$\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})=H_{1}({\mathbf{x}},{\mathbf{u}})+H_{2}({\mathbf{x}},{\mathbf{d}})+H_{3}({\mathbf{x}},{\mathbf{u}})H_{4}({\mathbf{x}},{\mathbf{d}}).$ (17)
Suppose that for each ${\mathbf{x}}\in{\mathcal{C}}$, there exists ${\mathbf{u}}\in\real^{m}$ such that:
  • there exist constants $a,b,a_{2},b_{2},a_{4},b_{4},b_{3}$ satisfying $a\leq\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})\leq b$, $a_{2}\leq H_{2}({\mathbf{x}},{\mathbf{d}})\leq b_{2}$, $a_{4}\leq H_{4}({\mathbf{x}},{\mathbf{d}})\leq b_{4}$, and $|H_{3}({\mathbf{x}},{\mathbf{u}})|\leq b_{3}$ almost surely;
  • there exist $\epsilon,\epsilon_{1},\epsilon_{2}>0$ such that $\epsilon_{1}+b_{3}\epsilon_{2}\leq\epsilon$, $\frac{\beta}{2}\geq 2\exp\{-\frac{2N\epsilon_{1}^{2}}{(b_{2}-a_{2})^{2}}\}$, $\frac{\beta}{2}\geq 2\exp\{-\frac{2N\epsilon_{2}^{2}}{(b_{4}-a_{4})^{2}}\}$, and
@@
Let ${\mathbf{x}}\in{\mathcal{C}}$ and pick ${\mathbf{u}}\in\real^{m}$ so that the conditions in the statement hold. Note that such $\mathbf{u}$} is dependent on the dataset ${\mathbf{D}}_{N}$. Now, we have
@@
$=\mathbb{P}_{N}\Big(\Big\rvert\tilde{H}_{2}({\mathbf{x}})+H_{3}({\mathbf{x}},{\mathbf{u}})\tilde{H}_{4}({\mathbf{x}})\Big\rvert\leq\epsilon\Big),$
where
$\tilde{H}_{2}({\mathbf{x}})=\mathbb{E}[\Delta H_{2}({\mathbf{x}},{\mathbf{d}})]-\frac{1}{N}\sum_{i=1}^{N}H_{2}({\mathbf{x}},{\mathbf{d}}^{(i)}),$
$\tilde{H}_{4}({\mathbf{x}})=\mathbb{E}[H_{4}({\mathbf{x}},{\mathbf{d}})]-\frac{1}{N}\sum_{i=1}^{N}H_{4}({\mathbf{x}},{\mathbf{d}}^{(i)}).$
Now, let $A_{H,2}$ be the event that $|\tilde{H}_{2}({\mathbf{x}})|\leq\epsilon_{1}$. Since $\frac{\beta}{2}\geq 2\exp\{-\frac{2N\epsilon_{1}^{2}}{(b_{2}-a_{2})^{2}}\}$, by Hoeffding’s inequality [39] we have that $\mathbb{P}_{N}(A_{H,2})\geq 1-\frac{\beta}{2}$. Similarly, let $A_{H,4}$ be the event that $|\tilde{H}_{4}({\mathbf{x}})|\leq\epsilon_{2}$. Since $\frac{\beta}{2}\geq 2\exp\{-\frac{2N\epsilon_{2}^{2}}{(b_{4}-a_{4})^{2}}\}$, by Hoeffding’s inequality [39] we have that $\mathbb{P}_{N}(A_{H,4})\geq 1-\frac{\beta}{2}$. Therefore, $\mathbb{P}_{N}(A_{H,2}\cap A_{H,4})\geq 1-\beta$. Now, note that if the event $A_{H,2}\cap A_{H,4}$ holds, since $\epsilon_{1}+b_{3}\epsilon_{2}\leq\epsilon$, we have that
$\Big\rvert\mathbb{E}[\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})]-\frac{1}{N}\sum_{i=1}^{N}\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}}^{(i)})\Big\rvert\leq\epsilon$
also holds. Therefore,
$\mathbb{P}_{N}\Big(\rvert\mathbb{E}[\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})]-\frac{1}{N}\sum_{i=1}^{N}\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}}^{(i)})\Big\rvert\leq\epsilon\Big)\geq 1-\beta.$
This, together with (18), implies that
@@
(Conditions of Proposition 7): The separable form of $\Delta h$ in (17) enables the use of Hoeffding’s inequality in the proof. Indeed, if (17) does not hold, since $\mathbf{u}$} is possibly dependent on ${\mathbf{D}}_{N}$, the quantities $\{\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}}^{(i)})\}_{i=1}^{N}$ are not necessarily independent and Hoeffding’s inequality can not be applied. On the other hand, a simple scenario in which $a,b$ as in Proposition 7 exist is if $h$ is uniformly bounded (which we can always assume without loss of generality). Moreover, $\frac{\beta}{2}\geq 2\exp\{-\frac{2N\epsilon_{1}^{2}}{(b_{2}-a_{2})^{2}}\}$, $\frac{\beta}{2}\geq 2\exp\{-\frac{2N\epsilon_{2}^{2}}{(b_{4}-a_{4})^{2}}\}$ hold if $N$ is sufficiently large. We also note that a version of Proposition 7 can be stated in which the constants defined therein are not uniform and instead depend on $\mathbf{x}$} or $\mathbf{u}$}. Although the assumption that such constants are uniform across $\mathbf{x}$} and $\mathbf{u}$} might add some conservatism, it leads to a simpler dependency on $\mathbf{u}$} in (18). In fact, (18) is convex in $\mathbf{u}$} under assumptions such as the ones in Remark 1. $\bullet$
@@
which is feasible by assumption. Now, by [40, Theorem 2.4], for each ${\mathbf{x}}\in{\mathcal{C}}$, $\mathbb{P}_{N}\Big(\mathbb{P}(\Delta h({\mathbf{x}},{\mathbf{u}},{\mathbf{d}})\leq 0)>\delta\Big)\leq\beta$. Equivalently,
@@ V-A Simulation
In simulation, we model the uncertainties in the terrain through a Gaussian disturbance with standard deviation $\sigma$ acting only in the $y$ direction, i.e., ${\mathbf{d}}_{k}=[0,d_{k},0]$, with $d_{k}\sim{\mathcal{N}}(0,\sigma)$. We consider a time horizon of 20 steps and take $\delta=0.1$, $\beta=0.01$, $\alpha=0.01$ and $\Delta t=0.1$. We consider a nominal controller ${\mathbf{k}}_{\text{nom}}({\mathbf{x}})=[0.2,0,-\theta]$. For each of the different conditions ——Markov (in the form of (8)), Cantelli, Hoeffding (note that the condition (17) holds in this example), and scenario approach——, we compute the control input at every state through a safety filter, i.e., by solving an optimization problem that finds the closest input to the nominal controller that satisfies the corresponding condition. We generate datasets of size 3252 and 113, for the Hoeffding and scenario approaches, respectively. For the Markov, Hoeffding and scenario approach-based condition, the optimization problem at every state is a convex quadratically constrained quadratic program (QCQP), which we solve using the cvxpy library in Python. The Cantelli-based condition leads to a non-convex problem, which we solve using the trust region solver of the minimize function in the SciPy library in Python. We implement each of the approaches (along with the nominal controller and the approach based on [27], which we refer to as the Martingale approach) in 400 different trajectories with initial condition ${\mathbf{x}}_{0}=[0,0,0]$.
@@
Figure 1: (left) Number of unsafe trajectories for different values of the disturbance variance $\sigma$. (right) Number of unsafe trajectories for different values of the trajectory length. No filter refers to the implementation of the nominal controller, whereas Markov, Cantelli, Hoeffding, and Scenario refer to the approaches in Propositions 278, respectively.
@@
(b) Markov
(c) Cantelli
(d) Hoeffding
(e) Scenario
Figure 2: Evolution of 400 different trajectories with initial condition at ${\mathbf{x}}_{0}=[0,0,0]$ and $\sigma=0.06$ ($\sigma=0.02$ for Markov) and trajectory length 2020 for different methods introduced in the paper, along with the nominal controller (left).
Markov Cantelli Hoeffding Scenario Martingale
$T_{ex}$ 3.47 3.78 3.89 79.96 3.46
$\sigma_{0}$ 0.03 0.16 0.13 0.21 0.25
TABLE II: Average time it takes to compute the controller ($T_{ex}$) in milliseconds and smallest value of $\sigma$ that leads to infeasibility $\sigma_{0}$ for the different approaches.
Finally, Table II shows the average execution times for the different approaches as well as the smallest value of $\sigma$ that leads to infeasibility of the corresponding condition.
The choice of which of the different approaches should be used depends on a variety of factors. Firstly, on what information is available about the uncertainty distribution. Secondly, on the desired level of conservativeness with respect to the safety constraints. Figure 1 and Table II show that there is a tradeoff between the empirical number of safety violations and the smallest variance that leads to infeasibility. Approaches yielding less safety violations generally lead to infeasible conditions for smaller values of $\sigma$. Thirdly, another point to consider is the computational time (cf. Table II). Although the martingale approach shows a larger value of $\sigma_{0}$ in Table II, it fails to meet the prescribed safety guarantees that the other approaches successfully satisfy.